OpenStudy (anonymous):

ƒ(x)=6x-5/x-3 List all horizontal asymptotes of ƒ. If there are no horizontal asymptotes, enter 'NONE'. Horizontal asymptotes y?

myininaya (myininaya):

Let f and g be polynomials \[\lim_{x \rightarrow \infty } \frac{f(x)}{g(x)}\] *If deg(f)=deg(g), then the horizontal asymptote is y= coefficient of the term with the biggest exponent on top/ coefficient of the term with the biggest exponent on bottom *If deg(f)>deg(g), then no horizzontal asymptote *If deg(f)<deg(g), then y=0 is the horizontal asymptote

myininaya (myininaya):

Where deg( ) means degree of

myininaya (myininaya):

So both your f and g are polynomials and what can you say about there degrees?

OpenStudy (anonymous):

Is this a trick ?

OpenStudy (anonymous):

so what's the answer? 3?

myininaya (myininaya):

No

myininaya (myininaya):

for your f what is deg(f)? for your g what is deg(g)? After you decide this, determine what if part above your problem satisfies to find the conclusion to your problem.

OpenStudy (anonymous):

Thanks so clear now, the answer is 6! hehe

OpenStudy (anonymous):

If the distance between the graph of a function and some fixed line approaches Zero as point on the graph moves increasingly far from the origin, we say that the graph approaches a line asymptotically and the line is an asymptote of the graph. For horizontal asymptote, a line \( y=b\) is horizontal asymptote of the graph of a function \( y=f(x) \) if either \( \lim \limits_ {x\to \infty } f(x)= b\) or \( \lim \limits_ {x\to -\infty } f(x)= b\)

OpenStudy (anonymous):

So, just take the limit to infinity, what Myin's gave you is simpler and faster way to detect the asymptote.

OpenStudy (anonymous):

Thanks FoolForMath!!

OpenStudy (anonymous):

Another cute thing to note is that "*If deg(f)>deg(g), then no horizzontal asymptote" but there will be oblique asymptote.